A steel wire fixed at both ends has a fundamental frequency of 200 Hz. A person can hear sound of maximum frequency 14 kHz. What is the highest harmonic that can be played on this string which is audible to the person?

Grade:10

1 Answers

Kevin Nash
9 years ago
Sol. Fundamental frequency f base 1 = 200 Hz Let l base 4 Hz be nth harmonic ⇒ F base2/F base 1 = 14000/200 ⇒ NF base1/F base1 = 70 ⇒ N = 70 ∴ The highest harmonic audible is 70^th harmonic.